### v2

@@ Abstract
Our survey is structured to provide an in-depth analysis of LLM applications across clinical settings, medical text data processing, research, education, and public health awareness. We begin by exploring the roles of LLMs in different medical applications, detailing how they are evaluated based on their performance in tasks such as clinical application, medical text data processing, information retrieval, data analysis, medical scientific writing, educational content generation etc. The subsequent sections delve into the methodologies employed in these evaluations, discussing the benchmarks and metrics used to assess the models’ effectiveness, accuracy, and ethical alignment.
@@ 2 Taxonomy and Structure of the Survey
  • In 3.2 Comprehensive Discussion on Evaluation Methods and Indicators, we will discuss the methodologies employed in evaluating LLMs, addressing questions about the nature of LLMs used (commercial vs. self-developed), the types of inputs (public datasets or proprietary data), the specifics of the prompts and tasks, and the benchmarks against which outputs are measured. This section aims to illuminate the depth and breadth of evaluation practices, highlighting the diversity of criteria from technical performance to ethical considerations.
@@ 3.1.2 Medical Text Data Processing Application Evaluations
In a study[60] on leveraging transformer-based models for structuring free-text radiological reports, the authors experiment with BERT, focusing on German chest X-ray reports from intensive care unit (ICU) patients. They evaluate the model’s performance on text classification tasks, specifically identifying six radiological findings mentioned in the reports. The evaluation metrics include macro-averaged F1-scores (MAF1) and confidence intervals (CIs). The study compares the efficacy of different pre-training and labeling strategies: an on-site pre-trained model using masked language modeling (Tmlm) versus a medically pre-trained model (Tmed), both fine-tuned with ’silver’ (rule-based) and ’gold’ (manually annotated) labels. The highest performance was observed for Tmlm,gold, achieving a MAF1 of 95.5% (CI: 94.5–96.3), highlighting the efficiency of custom pre-training combined with manual annotations for medical text data processing applications.

### v3

@@ Abstract
Our survey is structured to provide an in-depth analysis of LLM applications across clinical settings, medical text data processing, research, education, and public health awareness. We begin by exploring the roles of LLMs in various medical applications, detailing their evaluation based on performance in tasks such as clinical diagnosis, medical text data processing, information retrieval, data analysis, and educational content generation. The subsequent sections offer a comprehensive discussion on the evaluation methods and metrics employed, including models, evaluators, and comparative experiments. We further examine the benchmarks and datasets utilized in these evaluations, providing a categorized description of benchmarks for tasks like question answering, summarization, information extraction, bioinformatics, information retrieval and general comprehensive benchmarks. This structure ensures a thorough understanding of how LLMs are assessed for their effectiveness, accuracy, usability, and ethical alignment in the medical domain.
@@ 2 Taxonomy and Structure of the Survey
  • In 3.2 Comprehensive Discussion on Evaluation Methods and Metrics, we delve into the methodologies employed in evaluating LLMs, addressing key aspects such as the models used, evaluators, comparative experiments, and the diverse range of evaluation indicators.
    • In 3.2.1 Models, we discuss the various types of LLMs applied in medical evaluations. This includes commercial models like GPT-4, Claude, Bard, and others, as well as open-source models such as BERT, BioBERT, and PubMedBERT. The section also highlights customized models fine-tuned for specific medical tasks, offering insights into how different models are leveraged and assessed in the medical field.
    • 3.2.2 Evaluators and Comparative Experiments explores the diversity in evaluators, ranging from human experts to automated metrics and AI-driven assessments. We examine the roles of these evaluators in the context of different comparative experimental setups, which include comparisons between various LLMs, traditional NLP algorithms, and human experts. This subsection emphasizes the importance of rigorous and varied evaluation methodologies to ensure comprehensive assessment of LLMs’ performance.
    • In 3.2.3 Evaluation Metrics, we provide an extensive overview of the metrics used to evaluate LLM applications in the medical domain. These include correctness metrics (such as accuracy, precision), completeness metrics, composite metrics (such as F1-score and AUC), usability metrics (including helpfulness, safety, and robustness) and consistency & similarity metrics (including align with consensus and similarity). This section highlights how these metrics are applied across different medical tasks and scenarios to assess the efficacy, reliability, and safety of LLM outputs.
@@ 3.1.2 Medical Text Data Processing Application Evaluations
In a study[60] on leveraging transformer-based models for structuring free-text radiological reports, the authors experiment with BERT, focusing on German chest X-ray reports from intensive care unit (ICU) patients. They evaluate the model’s performance on text classification tasks, specifically identifying six radiological findings mentioned in the reports. The evaluation metrics include macro-averaged F1-scores (MAF1) and confidence intervals (CIs). The study compares the efficacy of different pre-training and labeling strategies: an on-site pre-trained model using masked language modeling ($T_{mlm}$) versus a medically pre-trained model ($T_{med}$), both fine-tuned with ’silver’ (rule-based) and ’gold’ (manually annotated) labels. The highest performance was observed for $T_{mlm}$,gold, achieving a MAF1 of 95.5% (CI: 94.5–96.3), highlighting the efficiency of custom pre-training combined with manual annotations for medical text data processing applications.